OpenStudy (anonymous):

Use Cramer's Rule to solve the system or to determine that the system is inconsistent or contains dependent equations. 2x + 4y = 12 3x + y = -2 1) inconsistent system 2) {(-2,4)} 3) {(-4, -2)} 4) {(4, -2)}

OpenStudy (ash2326):

@Audrae_World do you know cramer's rule?

OpenStudy (anonymous):

x + y + z = x – y – z = 0 x + 2y + z = 0

OpenStudy (ash2326):

That's not Cramer's rule Suppose I have the following system of equations \[a1x+b1y=c1\] and \[a2x+b2y=c2\] Then x and y are given as \[x=\frac{\left[\begin{matrix}c1 & a1 \\ c2& a2\end{matrix}\right]}{\left[\begin{matrix}a1 & b1 \\ a2& b2\end{matrix}\right]}\] \[y=\frac{\left[\begin{matrix}b1 & c1 \\ b2& c2\end{matrix}\right]}{\left[\begin{matrix}a1 & b1 \\ a2& b2\end{matrix}\right]}\] All are determinants, can you find x and y?

OpenStudy (anonymous):

WOAH my teacher didn't teach or classify with my class D:

OpenStudy (ash2326):

Do you know determinants ?

OpenStudy (anonymous):

determinants like this: [a b c d] and you multiply ad -bc?

OpenStudy (ash2326):

Yeah you're right, try this here

OpenStudy (anonymous):

Okay thanks :D

OpenStudy (ash2326):

Can you do it here? Did you understand?

OpenStudy (amistre64):

I think later on the cramer is refered to as the wronskian